# Automated Review Generation Method Based on Large Language Models

### v1

@@ 2.4 Hallucination mitigation
2.4.1 Prompt design and task decomposition
Firstly, we utilized strict and clear text summary guiding prompts, aimed at enhancing the scientific rationality of LLM’s outputs and ensuring accuracy and reliability in its analysis and generation processes. Notably, the task of automated review generation aligns well with the strengths of LLMs —information extraction and text generation capabilities. LLMs can rapidly and accurately extract core information from a vast array of literature and integrate it into a coherent and rigorous review text. To enhance efficiency and quality, we deconstructed the core of the review writing process, namely literature reading and summarization, into a series of text summarization tasks. This approach is adopted because summaries generated by LLM significantly surpass manually crafted and fine-tuned model-generated summaries in terms of fluency, factual consistency, and flexibility[46]. By establishing a list of questions, we directed the model to extract relevant content from the literature and respond based on this content, subsequently conducting a comprehensive analysis of all literature citations and responses. Ultimately, the LLM generates high-quality paragraphs closely related to the topic. Additionally, we employed a single-round, segmented generation strategy to avoid truncation limitations of approximately 8K output length. By reasonably segmenting long texts for generation, we not only ensured that the output was completed in a single conversational round but also provided finer parallel granularity to improve generation efficiency. In practice, we divided the 35 questions into 5 groups, ensuring that the generation results for each group could be successfully completed within the 8K limit of the LLM. This granularity avoids efficiency drops due to a high proportion of shared content and identical prompt frameworks, thereby enhancing processing speed while ensuring the quality of text generation.
2.4.2 Hallucination filtering and verification
To mitigate and rectify hallucinations, we employed a layered filtering and verification approach:
@@ 2.4.2 Hallucination filtering and verification
2.4.3 Effectiveness of hallucination mitigation
In evaluating the effectiveness of hallucination mitigation, we employed a confusion matrix to classify outcomes according to whether the LLM provided content and its pertinence to the original text, differentiating between two types of inaccuracies: false positives, which include fabricated or inconsistent information, and false negatives, referring to overlooked or partially extracted content. Our focus was primarily on reducing false positives, while adopting a relatively tolerant stance on false negatives.
Substantial progress was made in mitigating hallucinations. During paragraph generation, only 36% of outputs met criteria following format and DOI validations. This was achieved through 9 repetitions of generating 35 paragraphs, cumulatively resulting in 875 generations. Analyzing 343 relevant articles, we executed 1715 information extractions across 35 questions, yielding 8575 responses and ultimately aggregating to 2783 valid information combinations. Impressively, 84.80% of these outcomes were confirmed by the LLM as 100% consistent with the aggregated results (see Table 3 and Figure 2 (a)), affirming the model’s reliability. This method also establishes an rough benchmark for hallucination ratio, facilitating the selection and evaluation of LLMs.
Upon conducting manual verification on 25 articles each from the knowledge extraction and data mining stages, we calculated the accuracy, false positive rate, 95% confidence interval of the false positive rate, precision, recall, F1 score, and consistency (see Table 3). The 95% confidence interval for the false positive rate was provided by the statsmodels library in Python3. The results are detailed in the following table.
@@ 2.4.3 Effectiveness of hallucination mitigation
The data comparison underscores the efficacy of self-consistency verifications, revealing a substantial decrease in hallucinations, i.e., false positive content, while also compensating for some false negatives, where information was not fully extracted (see Figure 2 (b)). In the knowledge extraction phase, critical for review content, our manual sampling found no fabricated conclusions by LLMs (see Figure 2 (a)), attesting to our method’s scientific integrity and reliability. From the sampling results, we are over 95% confident that the likelihood of hallucinations in this part is less than 0.5% (see Table 3). Analysis of false positives in the post-aggregation data mining phase revealed hallucinations typically involved correct numerical extraction but with errors in units or definitions. False negatives mainly stemmed from LLMs’ inability to comprehend highly abstract expressions, reflecting a general LLM’s limited understanding of highly specialized scientific concepts. The incidence of hallucinations in knowledge extraction was significantly lower than in data mining, as answering questions did not involve converting units and concepts, thus avoiding the most challenging part of testing an LLM’s grasp of scientific knowledge. Domain-specific models enhanced by domain-adaptive pretraining (DAPT) [49]are poised to mitigate this issue. Opting not to fine-tune LLMs for specific domains in this study prioritizes out-of-the-box functionality and multi-domain generalization, utilizing a general LLM as the base. Comparisons between RAG and fine-tuning effects in specific domains indicate that RAG sustains efficacy with contextually new knowledge and offers a significantly lower initial cost[50], aligning with our objective to support researchers’ entry into diverse fields efficiently.
Considering the stringent accuracy requirements in research, increasing the number of repetitions can significantly reduce the probability of hallucinations appearing in aggregated results. Binomial probability calculations indicate that theoretically, a model with 79.09% accuracy yields aggregated prediction accuracies of 93.49%, 96.12%, and 97.64% after five, seven, and nine independent predictions, respectively, aligning with our sampling results (see Table 3). Detailed sampling outcomes and calculations are available in the SI.
@@
3 Discussion
In this study, we introduce an innovative LLM-based automated review generation method, adeptly addressing two key scientific challenges: streamlining literature review efficiency and significantly reducing LLM hallucination risks. This modular, comprehensive, end-to-end solution integrates modules for literature search, topic formulation, knowledge extraction, and review composition, transforming an extensive corpus of scientific literature into coherent, detailed, and error-free reviews tailored to specific research themes. Notably, our advanced data mining module offers experienced users an in-depth field overview, exploiting the LLM’s analytical prowess. Additionally, a user-friendly one-click program on Windows platforms significantly simplifies the review generation process.
A pivotal achievement of our method is its capacity to surpass traditional human resource limitations. Our rigorous quality assurance solution, encompassing format filtering, DOI verification, relevance verification and self-consistency verifications, ensures high reliability and traceability throughout the data processing pipeline. Expert evaluations with a case study of PDH catalysts confirm the method’s efficacy, with reviews paralleling manual ones in length and citations, but without hallucinations and with impeccable reference accuracy. Through rigorous testing, including the analysis of 875 LLM outputs from a sample of 25 articles, we demonstrate over 95% confidence in reducing the hallucination probability to below 0.5% (see Table 3).
Our method’s modular design offers excellent reusability and scalability. Individual modules like literature search, topic formulation, and knowledge extraction can serve various research purposes, like literature tracking, research topic discovery, and data mining datasets construction. Future enhancements will focus on augmenting LLM’s comprehension of scientific concepts through pan-scientific field fine-tuning, elevating the method’s overall utility. Planned upgrades include improving multimodal processing, automating scientific inquiries, personalizing text generation, and delving deeper into specific research areas.
In summary, our method signifies a major advancement in scientific research tools, offering rapid access to field breakthroughs and developments. It’s set to transform the landscape of scientific research, with far-reaching implications for knowledge base construction, literature recommendation, and structured academic writing, heralding a new era in scientific research productivity and interdisciplinary collaboration.
@@ 4 Methods
The method for constructing review articles consists of four parts: literature search, topic formulation, knowledge extraction and review composition, along with an additional data mining module for experienced users (see Figure 3).
Figure 3: Flowchart of the automated review generation method based on large language models. It includes four modules: i) literature search, ii) topic formulation, iii) knowledge extraction, iv) review composition, as well as an additional data mining module.
@@ 4.1 Literature search
Initially, a list of journals designated for the set review topic’s subject area is obtained from journal classification tables. Then, literature containing specified keywords within these selected journals is retrieved via search engine’s API. This is followed by a preliminary filter, checking each title and abstract for intersections with a selected list of keywords. Literature with intersections is saved, and those of a review nature are marked (see Figure 3 (i)). Our method supports various types of textual literature, including journals, patents, conference papers, books, etc. This means that any content in textual form can be included in the search scope, further expanding the application scenarios and coverage of our method. In our example, using "propane dehydrogenation" as a keyword, we retrieved 343 publications in top-tier Chemistry and Chemical Engineering journals (2022 Chinese Academy of Sciences classification), including 14 reviews, after filtering titles and abstracts with keywords like "propane dehydrogenation", "PDH", "ODH", "Oxidative Dehydrogenation", etc., through SerpAPI on Google Scholar.
@@ 4.2 Topic formulation
There are two approaches to constructing review topics: one involves LLM directly drafting the outline, and the other is based on LLM refining and drafting outlines from existing literature reviews. After obtaining a list of topics, additional topics can be manually added and sorted as needed (see Figure 3 (ii)). In our example, the Claude2 model generated an outline including 12 topics directly, and another with 9 topics and 35 guiding questions based on existing review articles (see Table 2).
@@ 4.3 Knowledge extraction
Based on the obtained list of topics, the LLM generates a list of questions for extracting information from literature, corresponding to each review topic. After repeating this process for multiple times for each article, all answers are concatenated. The LLM then determines whether the answers are relevant to the questions and aggregates them (see Figure 3 (iii)). In our example, in the case of PDH, after transforming the 35 guiding questions into questions for extracting information from literature, the Claude2 model was used to extract information from 343 top-tier articles five times, leading to the aggregation of 8575 responses into 2783 valid information combinations.
@@ 4.4 Review composition
After associating each article’s answers with their source DOI, paragraphs are generated and integrated for each topic. The LLM generates review paragraphs from all the answers combined, followed by summarization and outlook. After repeating multiple times, the LLM scores the generated paragraphs, selecting the best ones for each topic to form a preliminary draft of the full review. The full text is then polished with the help of the LLM, adjusting and checking citation formats to produce the final draft (see Figure 3 (iv)). In our example, each question’s answers from various articles were combined into JSON format information groups, inputted into the Claude2 model for paragraph generation, integrated to form smooth paragraphs, repeated 9 times, scored by the Claude2 model based on criteria (as shown in SI), and polished to produce the final draft.
@@ 4.5 Data mining
Based on the automated review generation method described above, we proposed a data mining method based on LLMs, catering to users with some domain knowledge. This method is almost identical to the knowledge extraction steps (see Figure 3 (ii)), effectively extracting and aggregating specific data from a large volume of literature. Users first define specific data extraction targets, which may include but are not limited to catalyst types, chemical compositions, and performance characteristics. On the established literature dataset, the LLM parses each article, extracting the user-defined target data multiple times and outputting in XML format. Similar to the knowledge extraction process (see Figure 3 (ii)), the LLM aggregates results from multiple extractions to finalize each article’s information for each extraction target. The extracted data often require manual cleaning and processing, including correcting extraction errors, standardizing data formats, and removing redundant information to facilitate subsequent statistical analyses. The cleaned data is then further integrated and analyzed to form visual charts. The code for the cleaning process and chart statistics can be generated by GPT4, requiring no programming background from the user. In our example, in the case of PDH catalysts, relevant literature from tiers one, two, and three was downloaded using the literature search module, filtered through abstracts and titles totaling 1041 articles, of which 839 were deemed PDH-related by the Claude2 model. After data cleaning and processing through Python3 programming, the extracted data included catalyst types, composition elements, active species elements, promoter elements, support materials, alloy structural types, alloy preparation methods, propane partial pressure, reaction temperature, inlet flow rate, selectivity, conversion, selectivity and other key indicators, covering seven categorical variables such as structure and element composition and three continuous variables related to reaction conditions. Subsequently, corresponding charts were generated through code execution. Initially, we tallied the annual publication numbers of various catalyst influencing factors, represented in line or Gantt charts. Furthermore, we calculated the average of the top five selectivity and stability for all influencing factors across all catalyst data, visualized in radar charts to showcase the peak performance achievable by a specific factor. Lastly, through pairwise combination analysis of influencing factors, we produced 45 bivariate correlation bubble charts, intuitively demonstrating how different variable combinations affect overall catalyst performance. These bubble charts use bubble color intensity, size, and border thickness to represent the levels of selectivity, conversion, and stability, respectively.
Data availability

### v3

@@ 2.5 Hallucination mitigation
Unlike search engines, LLMs’ process of understanding information and outputting it anew provides LLMs’ creativity while inevitably accompanying the "hallucination" phenomenon, referring to false information generated by LLMs without sufficient evidence support or contextually inconsistent, off-input information responses[40]. LLM hallucinations mainly originate from statistical biases in the training phase, noise in training data, and decision strategies when handling uncertain or multi-interpretable information during the alignment phase[40]. Currently, there is no solution within the field[9, 40, 41], and research suggests that hallucinations are unavoidable[42, 43]. Especially in specialized sub-fields, LLMs greatly exacerbate the hallucination phenomenon due to extremely scarce data exposure. For example, tests in literature[44] show that even the most advanced GPT-4 only has a 73.3% accuracy rate when answering professional multiple-choice questions, which is far from sufficient for scientific research fields that strictly require correctness. In scientific research, this may lead to unrealistic academic conclusions and misleading research directions, often meaning great waste of time and material consumption[41]. Therefore, effective hallucination mitigation is crucial for ensuring the scientific nature and reliability of automated review generation.
@@
Prompt design and task decomposition. Firstly, we utilized strict and clear text summary guiding prompts, aimed at enhancing the scientific rationality of LLM’s outputs and ensuring accuracy and reliability in its analysis and generation processes. Notably, the task of automated review generation aligns well with the strengths of LLMs—information extraction and text generation capabilities. LLMs can rapidly and accurately extract core information from a vast array of literature and integrate it into a coherent and rigorous review text. To enhance efficiency and quality, we deconstructed the core of the review writing process, namely literature reading and summarization, into a series of text summarization tasks. This approach is adopted because summaries generated by LLM significantly surpass manually crafted and fine-tuned model-generated summaries in terms of fluency, factual consistency, and flexibility[46]. By establishing a list of questions, we directed the model to extract relevant content from the literature and respond based on this content, subsequently conducting a comprehensive analysis of all literature citations and responses. Ultimately, the LLM generates high-quality paragraphs closely related to the topic. Additionally, we employed a single-round, segmented generation strategy to avoid truncation limitations of approximately 8K output length. By reasonably segmenting long texts for generation, we not only ensured that the output was completed in a single conversational round but also provided finer parallel granularity to improve generation efficiency. In practice, we divided the 35 questions into 5 groups, ensuring that the generation results for each group could be successfully completed within the 8K limit of the LLM. This granularity avoids efficiency drops due to a high proportion of shared content and identical prompt frameworks, thereby enhancing processing speed while ensuring the quality of text generation.
Hallucination filtering and verification. To mitigate and rectify hallucinations, we employed a layered filtering and verification approach:
@@
Effectiveness of hallucination mitigation. In evaluating the effectiveness of hallucination mitigation, we employed a confusion matrix to classify outcomes according to whether the LLM provided content and its pertinence to the original text, differentiating between two types of inaccuracies: false positives, which include fabricated or inconsistent information, and false negatives, referring to overlooked or partially extracted content. Our focus was primarily on reducing false positives, while adopting a relatively tolerant stance on false negatives.
Substantial progress was made in mitigating hallucinations. Specifically, in the paragraph generation part, which was achieved through 9 repetitions of 35-paragraph generation tasks, a total of 315 paragraphs that passed format checking and DOI checking were needed. Throughout the entire paragraph generation process, statistics show that LLM cumulatively performed 875 generations, of which only 36% of generation results passed after format and DOI list checks. In the analysis involving 343 topic-related literature, we divided 35 questions into 7 questions per segment for each literature, i.e., 5 segments per literature, totaling 1715 knowledge extractions. By conducting 5 repeated questions in each knowledge extraction, we obtained a total of 8575 answers, and finally aggregated 2783 effective information combinations after excluding answers unrelated to the literature and questions. Among these, up to 84.80% of the results were judged by LLM to have a 100% consistency rate when compared with the aggregated results (Table 1 and Fig. 4a), thus verifying the model’s stability. For specific methods, see the Methods section. This method also provides a rough standard for judging the proportion of hallucinations, which can be used in the screening and evaluation of LLMs.
To assess the effectiveness of the knowledge extraction and data mining stages, we implemented a rigorous manual verification process. Specifically, 25 randomly selected articles from each stage were evaluated by a third-year PhD student specializing in PDH research. For the knowledge extraction stage, 35 segments per article were examined, totaling 875 data points. The data mining stage assessed 14 catalyst properties, including 5 direct answer repetitions and final generated results, encompassing 1750 and 350 data points respectively. We employed precise classification criteria for the evaluation. In the knowledge extraction phase, true negatives (TN) were instances where the article did not address the guiding question and the LLM correctly identified it as irrelevant. True positives (TP) occurred when relevant topics were accurately extracted. False negatives (FN) were cases where relevant topics were incompletely extracted or incorrectly deemed irrelevant. False positives (FP) included irrelevant topics mistakenly identified as relevant or extractions that exceeded or deviated from the article’s actual content. Similar criteria were applied to the data mining stage, with particular attention to unit conversion errors, which were classified as false positives even if numerical values were correct. Consistency comparisons were conducted using the Claude2 model through designed prompt templates, comparing pre- and post-aggregation texts and statistically analyzing the model’s scoring results. Based on these evaluations, we calculated key metrics including accuracy, false positive rate (with 95% confidence intervals), precision, recall, F1 score, and consistency rate (Table 1). Confidence intervals for false positive rates were computed using Python3’s statsmodels library. For detailed results and calculation methods, see SI.
It is crucial to emphasize that this manual verification step was conducted to demonstrate the method’s effectiveness during the proof-of-concept phase and is not required in the actual automated review generation process. The detailed results are presented in Table 1.
@@
The data comparison underscores the efficacy of self-consistency verifications, revealing a substantial decrease in hallucinations, i.e., false positive content, while also compensating for some false negatives, where information was not fully extracted (Fig. 4b). In the knowledge extraction phase, critical for review content, our manual sampling found no fabricated conclusions by LLMs (Fig. 4a), attesting to our method’s scientific integrity and reliability. From the sampling results, we are over 95% confident that the likelihood of hallucinations in this part is less than 0.5% (Table 1), which is also the source of our confidence that this method supports fully automated processes without manual intervention. Analysis of false positives in the post-aggregation data mining phase revealed hallucinations typically involved correct numerical extraction but with errors in units or definitions. False negatives mainly stemmed from LLMs’ inability to comprehend highly abstract expressions, reflecting a general LLM’s limited understanding of highly specialized scientific concepts. The incidence of hallucinations in knowledge extraction was significantly lower than in data mining, as answering questions did not involve converting units and concepts, thus avoiding the most challenging part of testing an LLM’s grasp of scientific knowledge. Domain-specific models enhanced by domain-adaptive pretraining (DAPT)[49] are poised to mitigate this issue. Opting not to fine-tune LLMs for specific domains in this study prioritizes out-of-the-box functionality and multi-domain generalization, utilizing a general LLM as the base. Comparisons between RAG and fine-tuning effects in specific domains indicate that RAG sustains efficacy with contextually new knowledge and offers a significantly lower initial cost[50], aligning with our objective to support researchers’ entry into diverse fields efficiently.
Considering the stringent accuracy requirements in research, increasing the number of repetitions can significantly reduce the probability of hallucinations appearing in aggregated results. Binomial probability calculations indicate that theoretically, a model with 79.09% accuracy yields aggregated prediction accuracies of 93.49%, 96.12%, and 97.64% after five, seven, and nine independent predictions, respectively, aligning with our sampling results (Table 1). Detailed sampling outcomes and calculations are available in the SI. We believe that 5 repetitions is an ideal empirical value, and users do not need to change this parameter when using it.
@@
While both our method and RAG utilize LLM’s context learning ability, our approach fundamentally differs by achieving systematic knowledge reconstruction through multi-stage processing rather than simple retrieval combination. This method simulates the complete academic research process and produces coherent knowledge frameworks aligned with scholarly thinking, surpassing traditional RAG’s question-answering paradigm through comprehensive quality control and hallucination mitigation mechanisms.
3 CONCLUSIONS
In this study, we introduce an innovative LLM-based automated review generation method, addressing two fundamental challenges in scientific research: improving literature review efficiency and mitigating LLM hallucination risks. Through proposing an evaluation framework that ensures the objectivity and reliability via statistical validation, and innovatively compared LLM-generated reviews with high-quality manual reviews, we demonstrate that our modular end-to-end approach produces reviews comparable to or exceeding human-written ones, while maintaining high reliability and traceability. Expert evaluation using PDH catalysts as a case study confirms the method’s effectiveness: generated reviews are comparable to manual reviews in length and citations, show no hallucinations, and have impeccable reference accuracy. Statistical validation confirms the method’s effectiveness in hallucination reduction, with testing on 875 LLM outputs from 25 random articles showing hallucination probability below 0.5% at 95% confidence. The quality assurance pipeline ensures robust data processing. Additionally, our advanced data mining module offers experienced users’ in-depth field integrated perception, fully exploiting LLMs’ analytical capabilities. Furthermore, an open-source user-friendly one-click program developed for Windows platforms significantly simplifies the review generation process.
The method’s architecture offers significant advantages through its cross-disciplinary applicability without manual intervention or domain-specific knowledge injection. Its modular design enables component reuse for literature tracking, topic discovery, and dataset construction, while achieving cross-disciplinary generalization through LLM’s inherent contextual adaptability. This means that by providing corresponding domain literature input, the method can generate high-quality reviews across various disciplines. Future development will focus on enhanced multimodal processing capabilities, automated scientific question generation and answering, personalized text generation, integration with existing academic tools, and domain-specific features for structured data analysis.
This advancement heralds a new era in human-machine academic collaboration, offering broad prospects for LLMs as writing assistance tools. While not intended to replace traditional manual reviews, our method serves as a powerful auxiliary tool for rapid domain overview and research hotspot identification, laying the foundation for in-depth analysis. Beyond its demonstrated excellence in chemistry, the method’s technical framework exhibits remarkable cross-disciplinary applicability, potentially breaking down barriers between fields and catalyzing interdisciplinary innovation. By revolutionizing researcher-literature interaction and accelerating knowledge dissemination, this milestone advancement holds profound implications for knowledge base construction, literature recommendation, and structured academic writing, heralding a new era of scientific research productivity and interdisciplinary collaboration.
@@ 4 METHODS
Our method consists of four core components: literature search, topic formulation, knowledge extraction and review composition, along with a data mining module (Fig. 5a) and quality assessment framework (Fig. 5b). All prompt templates are available in SI and GitHub without requiring user adjustment.
Figure 5: a, Flowchart of the automated review generation method based on large language models. It includes four modules: i) literature search, ii) topic formulation, iii) knowledge extraction, iv) review composition, as well as an additional data mining module. b, Flowchart of the quality assessment framework for review generation based on large language models.
@@ 4.1 Literature search
Literature retrieval begins with journal selection from journal classification tables, followed by an API-based keyword search and preliminary title/abstract filtering using a keyword list, with review-type literature marked separately (Fig. 5a, i).
@@ 4.2 Topic formulation
Review topics can be constructed either through direct LLM outline drafting or through LLM extraction and refinement of existing literature reviews (Fig. 5a, ii). After obtaining a list of topics, additional topics can be manually added and sorted as needed. This manual addition is not mandatory but provides an interface for advanced users to intervene if necessary.
@@ 4.3 Knowledge extraction
Based on the topic list, LLM generates extraction questions and conducts multiple rounds of information retrieval from each article. The LLM evaluates answer relevance to questions using structured prompts, where combinations of questions, LLM-aggregated relevant answers, and their corresponding citations constitute valid information combinations for subsequent processing. For literature exceeding the context window of LLM, the text is segmented into approximately equal parts, processed separately, and results are integrated during answer aggregation (Fig. 5a, iii).
@@ 4.4 Review composition
Extracted answers are associated with source DOIs and integrated into topic-specific paragraphs. Through multiple iterations and LLM scoring, optimal versions are selected to form the preliminary draft, followed by citation verification and format standardization. For answers exceeding context window length, LLM performs compression based on referenced texts and extracted answers until fitting within window limits (Fig. 5a, iv).
@@ 4.5 Data mining
The data mining module extends knowledge extraction (Fig. 5a, iii) capabilities for specific data extraction and aggregation, enables extraction of user-defined targets (e.g., catalyst types, compositions, performance metrics) from literature. The LLM performs multiple rounds of parsing and extraction in XML format, followed by result aggregation. The extracted data undergoes standardization and cleaning, with GPT4-generated code facilitating statistical analysis and visualization, requiring no programming expertise from users.
4.6 Quality Assessment
The evaluation framework employs dual baselines using manual Q1 journal reviews and direct LLM generation for quality assessment (Fig. 5b). High-quality reviews are semantically segmented, with corresponding content regenerated using our method and compared against direct LLM-generated content. Assessment utilizes chain-of-thought prompts across 27 scoring items in 9 categories, implementing cross-evaluation and repetition strategies to mitigate bias. The page rerank algorithm converts relative comparisons to absolute scores on a 0-10 scale, with framework reliability validated through intraclass correlation coefficient (ICC) tests and transitive consistency ratio (TCR) analyses.
5 ACKNOWLEDGMENT
We acknowledge the Natural Science Foundation of China (No. 22121004), the Haihe Laboratory of Sustainable Chemical Transformations, the Program of Introducing Talents of Discipline to Universities (BP0618007) and the XPLORER PRIZE for financial support. We also acknowledge generous computing resources at High Performance Computing Center of Tianjin University.
6 AUTHOR CONTRIBUTIONS
J.G and Z.Z. conceived and supervised the project. S.W. and X.M. designed the research and developed the program. S.W. led the manuscript preparation. D.L. performed manual verification of hallucination. L.L., X.S., and X.C. advanced the integration of computational models, while X.L., R.L., C.P. and C.D. advanced the catalytic science. All authors contributed to writing and revising the manuscript.
7 DATA AVAILABILITY